OpenStudy (anonymous):

simplify the following: write in terms of a & b w/o logs or powers, where log2(3)=a log3(5)= b 1. log 2(3/2) 2. log3(45) 3. log2(1/9) 4. log16(3) 5. log4(27) I'm pretty sure I know how to do 1 and 3 1. log2(3/2)= a-1 3. log2(1/9)= -a^2 The rest i'm unsure...

OpenStudy (experimentx):

2) log3(45) = log3(9x5) = log3(3^3) + log3(5) = 3log3(3) + b = b+3

OpenStudy (experimentx):

4) let log16(3) = y you have => 3 = 16^y or 3 = 2^(4y) => log2(3)=4y or y = 1/4*a you can do the other in similar way

OpenStudy (anonymous):

the first answer is actually wrong because 3^3 = 27 so it's actually log3(3^2) + log3(5) => logb(b^x)=x so 2+b